In x dx/dy=2/xy^3
1
Expert's answer
2021-09-14T06:08:42-0400

Given "(ln x) \frac{dx}{dy}=\frac{2}{xy^3}"

It can be written as,


"(xln x){dx}=\frac{2}{y^3}dy"


"\frac{2}{y^3}dy = (xln x){dx}"


Integrating both sides,

"-\frac{1}{y^2}=[\frac{1}{2}x^2\ln \left(x\right)-\frac{x^2}{4}] + C"